# Movement Prediction-Adjusted Naive Forecast: Is the Naive Baseline Unbeatable in Financial Time Series Forecasting?

### v10

@@ Appendix B. Algebraic Derivation of $\Delta MSE_{\text{out}}$
$\quad-\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$
$=\frac{1}{M}\left(\sum_{t\in\text{correct}}\left(2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}-\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)-\sum_{t\in\text{incorrect}}\left(2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$ (B1)
@@
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx m_{\text{out}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx m_{\text{out}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out}}$ (B2)
@@
$\approx\frac{1}{M}\left(2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}\cdot(m_{\text{out}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out}})\right.-m_{\text{out}}^{\text{correct}}\cdot(\hat{\alpha}_{\text{in}}^{*})^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\left.-2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}\cdot(m_{\text{out}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out}})-m_{\text{out}}^{\text{incorrect}}\cdot(\hat{\alpha}_{\text{in}}^{*})^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)$
@@
Appendix C. Movement Prediction with an Exogenous Variable

### v11

@@ Appendix B. Algebraic Derivation of $\Delta MSE_{\text{out}}$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}|\epsilon_{t}|^{2}+\sum_{t\in\text{incorrect}}|\epsilon_{t}|^{2}\right)$
$\quad-\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(|\epsilon_{t}|^{2}-2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right$.
$\quad\left.+\sum_{t\in\text{incorrect}}\left(|\epsilon_{t}|^{2}+2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$
$=\frac{1}{N}\left(\sum_{t\in\text{correct}}\left(2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}-\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)-\sum_{t\in\text{incorrect}}\left(2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot|\epsilon_{t}|\cdot\bar{\epsilon}_{\text{in}}+\left(\hat{\alpha}_{\text{in}}^{*}\right)^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)\right)$ (B1)
@@
$\sum_{t\in\text{correct}}|\epsilon_{t}|\approx n_{\text{out}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out}},\quad\sum_{t\in\text{incorrect}}|\epsilon_{t}|\approx n_{\text{out}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out}}$ (B2)
@@
$\approx\frac{1}{N}\left(2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}\cdot(n_{\text{out}}^{\text{correct}}\cdot\bar{\epsilon}_{\text{out}})\right.-n_{\text{out}}^{\text{correct}}\cdot(\hat{\alpha}_{\text{in}}^{*})^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}$
$\quad\left.-2\cdot\hat{\alpha}_{\text{in}}^{*}\cdot\bar{\epsilon}_{\text{in}}\cdot(n_{\text{out}}^{\text{incorrect}}\cdot\bar{\epsilon}_{\text{out}})-n_{\text{out}}^{\text{incorrect}}\cdot(\hat{\alpha}_{\text{in}}^{*})^{2}\cdot\bar{\epsilon}_{\text{in}}^{2}\right)$
@@
Appendix C. Movement Prediction Using Co-moving Variables